# Nonequilibrium Probes of Quantum Geometry in Gapless Systems

### v1

@@ 4.2 Return probabilities for rotating drives
$f_{t}(\varphi)=f_{0}(\varphi)+\omega t$ (4.13)
where $\omega\equiv 2\pi/T>0$ is some frequency and $f_{0}\in\operatorname{Diff}(S^{1})$ is a given, fixed transformation. Any such family of deformations gives rise to a periodic Hamiltonian (2.13). Our goal is to find the resulting return probability (4.12) in the adiabatic limit $\omega L\to 0$, where $L$ is system size. We do this by first solving Eq. (4.8), which turns out to be integrable in the case of rotating drives. We then find the vector field $\mathbf{Y}_{t}$ in the group element (4.9), deduce the overlap $|\langle h|\,{\cal U}(f_{t}^{-1}g_{t})|h\rangle|^{2}$ at all times, and read off the return probability after one cycle.
@@
For rotating drives (4.13), the right-hand side of Eq. (4.8) is a vector field with profile $V_{t}(\varphi)=V_{0}(\varphi-\omega t)$, where $V_{0}(\varphi)=1/(f_{0}^{-1})^{\prime}(\varphi)$ owing to Eq. (4.6). It is a local velocity that rotates around the circle at constant angular velocity $\omega$. The corresponding flow $g_{t}$ is readily found with the same ansatz as in [66, 67]: denoting by $R_{\theta}(\varphi)\equiv\varphi+\theta$ the rotation by $\theta$, let
$g_{t}=R_{\omega t}\circ\zeta\circ R_{\Omega t}\circ\zeta^{-1}\circ g_{0},$ (4.14)
where the two unknowns are the deformation $\zeta\in\operatorname{Diff}(S^{1})$ and the angular velocity $\Omega$, both time-independent. The initial condition $g_{0}$ is arbitrary; choosing $g_{0}=f_{0}$ ensures that the initial state $|\psi(0)\rangle=\,{\cal U}(f_{0})|h\rangle$ is an eigenstate of the initial Hamiltonian, as in Sec. 4.1. Plugging the ansatz (4.14) into Eq. (4.8) and using the winding condition (2.2) yields
$(\zeta^{-1})^{\prime}(\varphi)=\frac{\Omega}{\frac{2\pi}{L}V_{0}(\varphi)-\omega},\qquad\frac{1}{\Omega}=\frac{1}{2\pi}\oint\frac{\textrm{d}\varphi}{\frac{2\pi}{L}V_{0}(\varphi)-\omega},$ (4.15)
which entirely determines the flow (4.14). [^11] This also shows that the ansatz (4.14) indeed is the solution of Eq. (4.8) for rotating drives. Note that this is only well-defined if $V_{0}(\varphi)-\frac{\omega L}{2\pi}$ never vanishes on the circle; in the adiabatic limit $\omega L\to 0$, this amounts to requiring that $V_{0}(\varphi)$ never vanishes, which is indeed satisfied by any velocity profile of the form (4.6).
Having found $g_{t}$, it is now straightforward to write the group element whose expectation value yields the return amplitude (4.10). With the initial condition $g_{0}=f_{0}$, Eq. (4.14) yields
$f_{0}^{-1}\circ g_{T}=f_{0}^{-1}\circ\zeta\circ R_{\Omega T}\circ\zeta^{-1}\circ f_{0}.$ (4.16)
This deformation is generally complicated—it is well away from the identity in $\operatorname{Diff}(S^{1})$—but we are specifically interested in the adiabatic regime $\omega L\ll 1$ where things simplify. To quickly see why, expand Eqs. (4.15) in the adiabatic limit and keep only the leading term, which yields $\Omega\sim 2\pi/L$ and $(\zeta^{-1})^{\prime}(\varphi)\sim 1/V_{0}(\varphi)=(f_{0}^{-1})^{\prime}(\varphi)$. Thus, $\zeta\sim f_{0}$ in the adiabatic limit, so Eq. (4.14) implies $\lim_{\omega L\to 0}g_{t}=f_{t}\circ R_{2\pi t/L}$ and the group element (4.16) boils down to a pure rotation $f_{0}^{-1}\circ g_{T}\sim R_{2\pi T/L}$. The latter gives rise to the expected dynamical phase in the amplitude (4.10). Our goal is to go beyond leading order by finding the small deviation of Eq. (4.16) away from a rotation.
@@
$\delta\equiv\frac{\omega}{\Omega}=\frac{1}{2\pi}\oint\frac{(L/T)\,\textrm{d}\varphi}{V_{0}(\varphi)-L/T}\ll 1,$ (4.17)
where $\Omega$ was defined in Eq. (4.15). This can be contrasted with the ‘bare’ adiabatic parameter $\omega L/2\pi=L/T$. The two are related by the adiabatic expansion of Eq. (4.15), namely
$\frac{1}{\delta}=\frac{T}{L}-\oint\frac{\textrm{d}\varphi}{2\pi f_{0}^{\prime}(\varphi)}-\frac{L}{T}\left[\oint\frac{\textrm{d}\varphi}{2\pi f_{0}^{\prime}(\varphi)^{2}}-\left(\oint\frac{\textrm{d}\varphi}{2\pi f_{0}^{\prime}(\varphi)}\right)^{2}\right]+O\left(\frac{L^{2}}{T^{2}}\right),$ (4.18)
confirming that $\delta\sim L/T$ to leading order. In practice, the first two terms of the expansion (4.18) often suffice. They are reminiscent of a dynamical phase and a Berry phase, to which they will be seen to correspond for $\mathrm{SL}(2,\mathbb{R})$ drives in Sec. 5.
It remains to compute the deformation $\zeta^{-1}\circ f_{0}$ of Eq. (4.16). The latter is close to the identity in the adiabatic limit: the expansion of Eq. (4.15) in powers of $L/T$ yields
$\zeta^{-1}(f_{0}(\varphi))=\varphi+\delta Z(\varphi)+O(\delta^{2}),\qquad Z(\varphi)\equiv\int_{0}^{\varphi}\textrm{d}\alpha\left(\frac{1}{f_{0}^{\prime}(\alpha)}-\oint\frac{\textrm{d}\theta}{2\pi f_{0}^{\prime}(\theta)}\right),$ (4.19)
where we used Eq. (4.18) to trade $L/T$ for $\delta$. As expected, this is an infinitesimal diffeomorphism involving a vector field $\delta Z(\varphi)\partial_{\varphi}$. One can now plug $\zeta^{-1}\circ f_{0}$ into Eq. (4.14) with $g_{0}=f_{0}$ to find
$f_{t}^{-1}(g_{t}(\varphi))=\varphi+\Omega t+\delta Y_{t}(\varphi)+O(\delta^{2})\qquad\text{with}\qquad Y_{t}(\varphi)\equiv Z(\varphi)-Z(\varphi+\Omega t).$ (4.20)
This is indeed a diffeomorphism of the form identified in Eq. (4.9), with a rotation by $\theta(t)=\Omega t$ and a small $O(\delta)$ correction involving the time-dependent vector field $\mathbf{Y}_{t}=Y_{t}(\varphi)\partial_{\varphi}$. The latter can be expressed in terms of $f_{0}(\varphi)$ thanks to Eq. (4.19). Note that this holds at all times; at the end of one cycle, one has $\Omega t=\Omega T=2\pi/\delta$ in terms of the adiabatic parameter (4.17).
@@
The return probability (4.12) is entirely determined by the Lie algebra element $\mathbf{Y}_{T}$, which we have just found in Eqs. (4.19)–(4.20). Using the quantum metric written in Eq. (2.25), we conclude that
@@
$\big|\langle\psi(0)|\psi(T)\rangle\big|^{2}\sim 1-\delta^{2}\oint\frac{\textrm{d}\varphi\,\textrm{d}\theta}{16\pi^{2}\sin^{2}([\varphi-\theta]/2)}\Biggl(\Bigl(h-\frac{c}{24}\Bigr)\Biggl[\int_{\theta}^{\theta+\frac{2\pi}{\delta}}\!\!\!\frac{\textrm{d}\alpha}{f_{0}^{\prime}(\alpha)}-\int_{\varphi}^{\varphi+\frac{2\pi}{\delta}}\!\!\!\frac{\textrm{d}\alpha}{f_{0}^{\prime}(\alpha)}\Biggr]^{2}\\ +\frac{c}{24}\Biggl[\frac{1}{f_{0}^{\prime}(\theta+\frac{2\pi}{\delta})}-\frac{1}{f_{0}^{\prime}(\theta)}-\frac{1}{f_{0}^{\prime}(\varphi+\frac{2\pi}{\delta})}+\frac{1}{f_{0}^{\prime}(\varphi)}\Biggr]^{2}\,\Biggr)$ (4.21)
@@
for adiabatic rotating drives in CFT, where we used $\Omega T=2\pi/\delta$. This is plotted in Fig. 1 as a function of $T/L\sim 1/\delta$ for different choices of the initial deformation $f_{0}$.
@@
A striking aspect of the plots in Fig. 1 is the oscillating convergence of the return probability in the adiabatic limit $T/L\to\infty$. These oscillations stem from the ubiquitous appearance, in Eq. (4.21), of the shift by $2\pi/\delta\sim T/L$ in the $2\pi$-periodic function $1/f_{0}^{\prime}(\varphi)$: depending on the value of $2\pi/\delta$, the shift gives rise to constructive or destructive ‘interferences’ of $1/f_{0}^{\prime}(\varphi)$. A more direct way to predict such oscillations is to refer to the expression (2.22a) of the Virasoro quantum metric in terms of Fourier modes. In the case at hand, the Fourier modes of $Y_{t}(\varphi)$ in Eq. (4.19) are $\hat{Y}_{m}(t)=\hat{Z}_{m}(1-e^{im\Omega t})$, where $\hat{Z}_{m}$ is the $m^{\text{th}}$ mode of $1/f_{0}^{\prime}(\varphi)$ (for $m\neq 0$). The exponential $e^{im\Omega t}$ leads to an oscillating pattern of the squared quantum norm
${\cal G}(\mathbf{Y}_{t},\mathbf{Y}_{t})=4\sum_{m\in\mathbb{Z}}|m|\Big(h+\frac{c}{24}(m^{2}-1)\Big)|\hat{Z}_{m}|^{2}\sin^{2}(m\Omega t/2)$ (4.22)
@@
as a function of $t$, reflecting the micromotion of $g_{t}$ around $f_{t}$ visible in Fig. 4. This micromotion is manifest in the time-dependent overlap $|\langle h|\,{\cal U}(f_{t}^{-1}\circ g_{t})|h\rangle|^{2}\sim 1-\delta^{2}{\cal G}(\mathbf{Y}_{t},\mathbf{Y}_{t})$ between the exact state $\propto\,{\cal U}(g_{t})|h\rangle$ and its adiabatic approximation $\propto\,{\cal U}(f_{t}\circ R_{2\pi t/L})|h\rangle$.
5 Rotating drives in $\mathrm{SL}(2,\mathbb{R})$

### v2

@@ IV.2 Return probabilities for rotating drives
$f_{t}(\varphi)=f_{0}(\varphi)+\omega t$ (55)
for some frequency $\omega\equiv 2\pi/T>0$ and a fixed transformation $f_{0}\in\operatorname{Diff}(S^{1})$. Any such transformations give rise to a periodic Hamiltonian (19). Our goal is to find the resulting return probability (54) in the adiabatic limit $\omega L\to 0$, where $L$ is system size. We do this by first solving Eq. (51), which turns out to be integrable in the case of rotating drives. We then find the vector field $\mathbf{Y}_{t}$ in the group element (52), deduce the overlap $|\langle h|\,{\cal U}(f_{t}^{-1}g_{t})|h\rangle|^{2}$ at all times, and read off the return probability after one cycle.
@@
For a rotating drive (55), the right-hand side of Eq. (51) is a vector field with profile $V_{t}(\varphi)=V_{0}(\varphi-\omega t)$, where $V_{0}(\varphi)=1/(f_{0}^{-1})^{\prime}(\varphi)$ owing to Eq. (49). It is a local velocity that rotates around the circle at constant frequency $\omega$. Let again $R_{\theta}(\varphi)\equiv\varphi+\theta$ denote rotation by $\theta$. Then the flow $g_{t}$ corresponding to a rotating drive is readily found with the same ansatz as in [97, 98]. Namely,
$g_{t}=R_{\omega t}\circ\zeta\circ R_{\Omega t}\circ\zeta^{-1}\circ g_{0}$ (56)
for a deformation $\zeta\in\operatorname{Diff}(S^{1})$ and a frequency $\Omega$, both of which are time-independent unknowns. The initial condition $g_{0}$ is arbitrary; choosing $g_{0}=f_{0}$ ensures that the initial state $|\psi(0)\rangle=\,{\cal U}(f_{0})|h\rangle$ is an eigenstate of the initial Hamiltonian, as in Sec. IV.1. Plugging the ansatz (56) into Eq. (51) and using the winding condition (8) yields
$(\zeta^{-1})^{\prime}(\varphi)=\frac{\Omega}{\frac{2\pi}{L}V_{0}(\varphi)-\omega},\quad\frac{1}{\Omega}=\oint\frac{\textrm{d}\varphi/2\pi}{\frac{2\pi}{L}V_{0}(\varphi)-\omega},$ (57)
which entirely determines the flow (56). [^15] Since $V_{t}(\varphi)$ given by Eq. (49) is continuous in $t$ and smooth in $\varphi$, the existence and uniqueness theorem ensures that the ansatz (56) is the solution of Eq. (51). Note that this is well defined only if $V_{0}(\varphi)-\frac{\omega L}{2\pi}$ never vanishes on the circle; in the adiabatic limit $\omega L\to 0$, this amounts to requiring that $V_{0}(\varphi)\neq 0$, which is indeed satisfied by any velocity profile of the form (49).
Having found $g_{t}$, it is now straightforward to write the group element whose expectation value yields the return amplitude (53). With the initial condition $g_{0}=f_{0}$, Eq. (56) yields
$f_{0}^{-1}\circ g_{T}=f_{0}^{-1}\circ\zeta\circ R_{\Omega T}\circ\zeta^{-1}\circ f_{0}.$ (58)
This deformation is generally complicated—it is well away from the identity in $\operatorname{Diff}(S^{1})$—but we are specifically interested in the adiabatic regime $\omega L\ll 1$ where things simplify. To quickly see why, expand Eqs. (57) in the adiabatic limit and keep only the leading terms, which yields $\Omega\sim 2\pi/L$ and $(\zeta^{-1})^{\prime}(\varphi)\sim 1/V_{0}(\varphi)=(f_{0}^{-1})^{\prime}(\varphi)$. Thus, $\zeta\sim f_{0}$ in the adiabatic limit, so Eq. (56) implies $\lim_{\omega L\to 0}g_{t}=f_{t}\circ R_{2\pi t/L}$ and the group element (58) boils down to a pure rotation $f_{0}^{-1}\circ g_{T}\sim R_{2\pi T/L}$. The latter gives rise to the expected dynamical phase in the amplitude (53). Our goal is to go beyond leading order by finding the small deviation of Eq. (58) away from a rotation.
@@
$\delta\equiv\frac{\omega}{\Omega}=\frac{1}{2\pi}\oint\frac{(L/T)\,\textrm{d}\varphi}{V_{0}(\varphi)-L/T}\ll 1,$ (59)
where $\Omega$ was introduced in Eq. (57). This can be contrasted with the ‘bare’ adiabatic parameter $\omega L/2\pi=L/T$. The two are related by the adiabatic expansion of Eq. (57), namely
$\frac{1}{\delta}=\frac{T}{L}-\oint\frac{\textrm{d}\varphi}{2\pi f_{0}^{\prime}(\varphi)}$ (60)
$\;\;\;-\frac{L}{T}\bigg[\oint\frac{\textrm{d}\varphi}{2\pi f_{0}^{\prime}(\varphi)^{2}}-\left(\oint\frac{\textrm{d}\varphi}{2\pi f_{0}^{\prime}(\varphi)}\right)^{2}\,\bigg]+O\Bigl(\frac{L^{2}}{T^{2}}\Bigr),$
confirming that $\delta\sim L/T$ to leading order. [We used $V_{0}(\varphi)=1/(f_{0}^{-1})^{\prime}(\varphi)$ from Eq. (49).] In practice, the first two terms of the expansion (60) often suffice. They are reminiscent of a dynamical phase and a Berry phase, to which they will be seen to correspond for $\mathrm{SL}(2,\mathbb{R})$ drives in Sec. V.
It remains to compute the deformation $\zeta^{-1}\circ f_{0}$ of Eq. (58). The latter is close to the identity in the adiabatic limit: the expansion of Eq. (57) in powers of $L/T$ yields
$\zeta^{-1}(f_{0}(\varphi))=\varphi+\delta W(\varphi)+O(\delta^{2}),$ (61)
$W(\varphi)\equiv\int_{0}^{\varphi}\textrm{d}\alpha\left(\frac{1}{f_{0}^{\prime}(\alpha)}-\oint\frac{\textrm{d}\theta}{2\pi f_{0}^{\prime}(\theta)}\right),$
where we used Eq. (60) to trade $L/T$ for $\delta$. As expected, this is an infinitesimal diffeomorphism involving a vector field $\delta W(\varphi)\partial_{\varphi}$. [^16] This $\mathbf{W}$ has nothing to do with the one used for linear response in Sec. III.2. One can now plug $\zeta^{-1}\circ f_{0}$ into Eq. (56) with $g_{0}=f_{0}$ to find
$f_{t}^{-1}(g_{t}(\varphi))$ $=\varphi+\Omega t+\delta Y_{t}(\varphi)+O(\delta^{2}),$ (62)
$Y_{t}(\varphi)$ $\equiv W(\varphi)-W(\varphi+\Omega t).$
This is indeed a diffeomorphism of the form identified in Eq. (52), with a rotation by $\theta(t)=\Omega t$ and a small $O(\delta)$ correction involving the time-dependent vector field $\mathbf{Y}_{t}=Y_{t}(\varphi)\partial_{\varphi}$. The latter can be expressed using $f_{0}(\varphi)$ thanks to Eq. (61). Note that this holds at all times; at the end of one cycle, one has $\Omega t=\Omega T=2\pi/\delta$ in terms of the adiabatic parameter (59).
@@
The probability (54) is entirely determined by the Lie algebra element $\mathbf{Y}_{T}$, which we just found in Eqs. (61)–(62). Using the quantum metric written in Eq. (32), we conclude that
@@
$\big|\langle\psi(0)|\psi(T)\rangle\big|^{2}\sim 1-\delta^{2}\oint\frac{\textrm{d}\varphi\,\textrm{d}\theta}{16\pi^{2}\sin^{2}([\varphi-\theta]/2)}$
$\times\Biggl[\Bigl(h-\frac{c}{24}\Bigr)\biggl(\int_{\theta}^{\theta+\frac{2\pi}{\delta}}\!\!\!\frac{\textrm{d}\alpha}{f_{0}^{\prime}(\alpha)}-\int_{\varphi}^{\varphi+\frac{2\pi}{\delta}}\!\!\!\frac{\textrm{d}\alpha}{f_{0}^{\prime}(\alpha)}\biggr)^{2}$
$\;\;\;+\frac{c}{24}\biggl(\frac{f_{0}^{\prime}(\theta){-}f_{0}^{\prime}(\theta+\frac{2\pi}{\delta})}{f_{0}^{\prime}(\theta)f_{0}^{\prime}(\theta+\frac{2\pi}{\delta})}-\frac{f_{0}^{\prime}(\varphi){-}f_{0}^{\prime}(\varphi+\frac{2\pi}{\delta})}{f_{0}^{\prime}(\varphi)f_{0}^{\prime}(\varphi+\frac{2\pi}{\delta})}\biggr)^{2}\Biggr]$ (63)
@@
for adiabatic rotating drives in CFT, where we used $\Omega T=2\pi/\delta$. This is plotted in Fig. 1 as a function of $T/L\sim 1/\delta$ for different initial deformations $f_{0}$.
@@
A striking aspect of the plots in Fig. 1 is the oscillating convergence of the return probability in the adiabatic limit $T/L\to\infty$. These oscillations stem from the ubiquitous appearance, in Eq. (63), of the shift by $2\pi/\delta\sim T/L$ in the $2\pi$-periodic function $1/f_{0}^{\prime}(\varphi)$: depending on the value of $2\pi/\delta$, the shift gives rise to constructive or destructive ‘interferences’ of $1/f_{0}^{\prime}(\varphi)$. A more direct way to predict such oscillations is to refer to the expression (29a) of the Virasoro quantum metric in terms of Fourier modes. In the case at hand, the Fourier modes of $Y_{t}(\varphi)$ in Eq. (61) are $\hat{Y}_{m}(t)=\hat{W}_{m}(1-e^{im\Omega t})$, where $\hat{W}_{m}$ is the $m^{\text{th}}$ mode of $1/f_{0}^{\prime}(\varphi)$ (for $m\neq 0$). The exponential $e^{im\Omega t}$ leads to an oscillating pattern of the squared quantum norm
${\cal G}(\mathbf{Y}_{t},\mathbf{Y}_{t})=4\sum_{m\in\mathbb{Z}}|m|\Big(h+\frac{c}{24}(m^{2}-1)\Big)\\ \times|\hat{W}_{m}|^{2}\sin^{2}(m\Omega t/2)$ (64)
@@
as a function of $t$, reflecting the micromotion of $g_{t}$ around $f_{t}$ seen in Fig. 4. This micromotion is manifest in the time-dependent overlap $|\langle h|\,{\cal U}(f_{t}^{-1}\circ g_{t})|h\rangle|^{2}\sim 1-\delta^{2}{\cal G}(\mathbf{Y}_{t},\mathbf{Y}_{t})$ between the exact state $\propto\,{\cal U}(g_{t})|h\rangle$ and its adiabatic approximation $\propto\,{\cal U}(f_{t}\circ R_{2\pi t/L})|h\rangle$.
Adding left movers.
The chiral case is readily generalized to a nonchiral CFT that involves both right and left movers, as discussed around Eq. (7) and in Appendix A. This is crucial for the effective description of lattice models and for the comparison with their numerics, such as those in Fig. 1.
The Hamiltonian of a driven nonchiral inhomogeneous CFT is a sum
$H_{\text{nonchiral}}(t)=\int\textrm{d}x\,v_{t}(x)\big[{\mathcal{T}}(x)+\overline{{\mathcal{T}}}(x)\big]$ (65)
of the right-moving term (1) with its left-moving partner. The same velocity profile $v_{t}(x)$ multiplies both light-cone components of the stress-energy tensor, namely ${\mathcal{T}}(x)=(2\pi/L)^{2}{\mathcal{T}}(\varphi)$ and $\overline{{\mathcal{T}}}(x)=(2\pi/L)^{2}\overline{{\mathcal{T}}}(\varphi)$ for $\varphi=2\pi x/L$. [^17] One could use different velocity profiles for right and left movers. This is more complicated to implement in lattice models, so we do not discuss such generalizations. While the Fourier expansion of $\mathcal{T}$} is provided by Eq. (13) in terms of right-moving Virasoro generators $L_{m}$, that of $\overline{{\mathcal{T}}}$ is
$\overline{{\mathcal{T}}}(\varphi)=\frac{1}{2\pi}\sum_{n\in\mathbb{Z}}e^{-in\varphi}\left(\bar{L}_{n}-\frac{\bar{c}}{24}\delta_{n,0}\right)$ (66)
in terms of left-moving Virasoro generators $\bar{L}_{m}$. The latter commute with the $L_{m}$s and satisfy the same commutation relations (15) of the Virasoro algebra, but with a left-moving central charge $\bar{c}$ that may differ from $c$. The nonchiral Hilbert space is a sum of tensor products of right and left representations of the Virasoro algebra, respectively generated by the $L_{m}$ and $\bar{L}_{m}$ operators. Nonchiral highest-weight states are tensor products of the chiral ones in Eq. (22).
The fact that right and left sectors commute means that the full return amplitude (44) in a nonchiral CFT is the product of right- and left-moving contributions that can be computed separately. The same is true of the return probability (54). In the same adiabatic expansion as before, the latter involves again the quantum metric, which is now a sum of contributions from right and left movers:
$\big|\langle\psi(0)|\psi(T)\rangle\big|^{2}\sim 1-\delta^{2}{\cal G}(\mathbf{Y}_{T},\mathbf{Y}_{T})-\bar{\delta}^{2}\bar{{\cal G}}(\bar{\mathbf{Y}}_{T},\bar{\mathbf{Y}}_{T}).$ (67)
Here all barred quantities denote left-moving analogues of (and generally differ from) the right-moving ones, and the state $|\psi(t)\rangle$ is that of the full nonchiral theory. In particular, $\bar{{\cal G}}$ has formulas analogous to Eqs. (29a)–(30a) or (32) with $c$ and $h$ replaced by $\bar{c}$ and $\bar{h}$, where $\bar{h}$ is the left-moving highest weight. The fact that the same velocity profile multiplies both $\mathcal{T}$} and $\overline{{\mathcal{T}}}$ in the driven Hamiltonian (65) further means that the left-moving flow $\bar{g}_{t}(\varphi)$ satisfies an almost identical equation to Eq. (51) for the right-moving flow $g_{t}(\varphi)$, the only difference being a sign:
$\partial_{t}\bar{g}_{t}\circ\bar{g}_{t}^{-1}=-\frac{2\pi}{L}V_{t}.$ (68)
The sign difference stems from the opposite directions of the flows, echoed by sign differences in many other formulas between right and left movers (see Appendix A). This includes the different conventions for the Fourier expansions (13) and (66), ensuring that the $L_{m}$s and $\bar{L}_{m}$s still satisfy the same commutations relations (15).
For rotating drives where $V_{t}(\varphi)=V_{0}(\varphi-\omega t)$, the fact that left movers experience a velocity opposite to that of right movers has important implications for the return probability (67). An analogue of Eq. (56) still holds, namely $\bar{g}_{t}=R_{\omega t}\circ\bar{\zeta}\circ R_{-\overline{\Omega}t}\circ\bar{\zeta}^{-1}\circ\bar{g}_{0}$, but solving the left-moving flow equation (68) now yields
$(\bar{\zeta}^{-1})^{\prime}(\varphi)=\frac{\overline{\Omega}}{\frac{2\pi}{L}V_{0}(\varphi)+\omega},\quad\frac{1}{\overline{\Omega}}=\oint\frac{\textrm{d}\varphi/2\pi}{\frac{2\pi}{L}V_{0}(\varphi)+\omega}.$ (69)
These only differ from the right-moving formulas in Eq. (57) by the sign in front of $\omega$. In particular, the left-moving adiabatic parameter is $\bar{\delta}\equiv\omega/\overline{\Omega}$, which differs from Eq. (59) by a sign in its denominator, and the Lie algebra element $\bar{\mathbf{Y}}_{T}$ for rotating drives can be found just as we did for $\mathbf{Y}_{T}$ in Eqs. (61)–(62). The return probability (67) is given by the product of Eq. (63) with its left-moving cousin, obtained by replacing $\delta$, $c$, $h$ by $\bar{\delta}$, $\bar{c}$, $\bar{h}$. Equivalently, it can be Fourier expanded in a way analogous to Eq. (64):
$\big|\langle\psi(0)|\psi(T)\rangle\big|^{2}\sim 1$ (70)
$-4\delta^{2}\sum_{m\in\mathbb{Z}}|m|\Bigr(h+\frac{c}{24}(m^{2}{-}1)\Bigl)|\hat{W}_{m}|^{2}\sin^{2}\Bigl(\frac{m\Omega T}{2}\Bigr)$
$-4\bar{\delta}^{2}\sum_{m\in\mathbb{Z}}|m|\Bigl(\bar{h}+\frac{\bar{c}}{24}(m^{2}{-}1)\Bigr)|\hat{W}_{m}|^{2}\sin^{2}\Bigl(\frac{m\overline{\Omega}T}{2}\Bigr).$
The CFT used in Fig. 1 has $c=\bar{c}=1$, and the initial state $|\psi(0)\rangle$ is the (deformed) nonchiral ground state with $h=\bar{h}=0$.
Note an important difference between the nonchiral formula (70) and its right-moving simplification (64). The latter vanishes at $t=T$ when $\Omega T$ is an integer multiple of $2\pi$, so the chiral return probability (63) is guaranteed to reach the value $|\langle\psi(0)|\psi(T)\rangle|^{2}=1$ repeatedly when $T$ tends to infinity. By contrast, the nonchiral probability (70) never quite reaches that value because $\Omega\neq\overline{\Omega}$, even though $\Omega\sim 2\pi/L\sim\overline{\Omega}$ in the adiabatic limit. The consequence of this slight mismatch between $\Omega$ and $\overline{\Omega}$ is visible e.g. in Fig. 1(b), where the maxima of the return probability all sit slightly yet strictly below $1$.
V Rotating drives in $\mathrm{\bf SL}\bm{(}\bm{2}\bm{,}\mathbb{R}\bm{)}$